A bar 40cm long with insulated sides had its ends A and B maintained at 0 degree and 100 degree Until steady state condition prevails Then the temperature at A is suddenly raised to 20 degree and at the same time that of B is lowered to 60 degree. Find the temperature distribution in the bar at time t
1
Expert's answer
2020-11-25T19:33:52-0500

The one dimensional heat flow equation is given by:

"\frac{\partial u}{\partial t}=\alpha^2\frac{\partial^2 u}{\partial x^2}"

In steady-state:

"\frac{\partial u}{\partial t}=0"

So:

"\frac{\partial^2 u}{\partial x^2}=0"

Solving:

u=ax+b

Initial conditions in steady-state:

u=0 at x=0

u=100 at x=40

Then:

u=2.5x

Hence the boundary conditions are:

u(0,t)=20

u(40,t)=60

u(x,0)=2.5x

The solution of one dimensional heat flow equation is given by:

u(x,t)=(A*cos("\lambda"x)+B*sin("\lambda"x))*exp["-\alpha^2*\lambda^2*t" ]

Break up the required funciton u (x,t) into two parts:

u(x,t)=us(x)+ut(x,t)

To find us(x):

"\frac{\partial^2 u_s}{\partial x^2}=0"

us(x) = ax+b

us(0)=20

us(40)=60

Then:

us=x+20

To find ut(x,t):

ut( x,t) = u(x,t) –us(x)

ut(0,t) = u(0,t)–us(0) = 20–20 = 0

ut(40,t) = u(40,t)–us(40) = 60–60 = 0

ut(x,0) = u(x,0) –us(x) = 2.5x-x-20=1.5x-20

ut(x,t)=(A*cos("\lambda"x)+B*sin("\lambda"x))*exp["-\alpha^2*\lambda^2*t" ]

A=0; "\lambda" =(n"\pi")/40

ut(x,t)=B*sin((n"\pi"x)/40)*exp["-\alpha^2*((n\pi)/40)^2*t" ]

Most general solution:

ut(x,t)="\displaystyle\sum_{n=1}^\infty" Bn*sin((n"\pi"x)/40)*exp["-\alpha^2*((n\pi)/40)^2*t" ]

Using condition:

Bn = 2/40*"\int_{0}^{40}" (1.5x-20)*sin(n"\pi"x/40)dx =


u=us+ut